OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} {n^2}/{2^n}\] I'm supposed to use the ratio test to see if this series is absolutely convergent, conditionally convergent, or divergent.

OpenStudy (anonymous):

\[{n+1}^{2}/ {2}^{n+1}\] is greater than the original equation

OpenStudy (turingtest):

you need the absolute value of the ratio of successive terms though

OpenStudy (anonymous):

sorry i meant (n+1)

OpenStudy (anonymous):

ohh i need to see if the absolute value of the ratio is greater or less than 1?

OpenStudy (experimentx):

\[ \lim_{n \rightarrow \infty} \left ( \frac{(n+1)^2}{2^{n+1}} \over \frac{n^2}{2^{n}}\right ) = \lim_{n \rightarrow \infty} \frac{ \left( \frac{n + 1}{n} \right )^2}{2} = \frac 1 2 < 1\]

OpenStudy (anonymous):

\[\frac{(n+1)^2\times 2^n}{2^{n+1}n^2}\]

OpenStudy (anonymous):

you get \[\frac{(n+1)^2}{2n^2}\] take limit as \(n\to \infty\) and get \(\frac{1}{2}\)

OpenStudy (anonymous):

makes sense...so this series is absolutely convergent

OpenStudy (anonymous):

if it is convergent it is certainly absolutely convergent because all the terms are positive

OpenStudy (anonymous):

btw experimentx is of course right, but i find it easier not to write a compound fraction if i can get away with it. i.e if your term \(a_n\) is a fraction multiply by the reciprocal when you take your ratio

OpenStudy (anonymous):

thanks so much everyone!!!

OpenStudy (anonymous):

whatever you are most comfortable with

OpenStudy (anonymous):